The apparent weight of a body holy emerged in the water is 32N and eat sweet in a is 96N Calculate the volume of the body. (Density of water =1000kgm-3,g=10ms-2
1
Expert's answer
2021-10-03T16:17:03-0400

The apparent weight of a body holy emerged in the water is 32N and eat sweet in a is 96N Calculate the volume of the body. (Density of water =1000kgm-3,g=10ms-2


The weight of the body in the water is the resultant of two forces: weight of the body in air "F_g = 96N" acting down and buoyant force "F_b" acting up:


"W = F_g - F_b"

The buoyant force is given as follows:

"F_b = \rho_w V g"

where "\rho_w = 1000kg/m^3" is the density of water and "V" is the volume of the body. Expressing "V" and substituting "W = 32N", obtain:


"V = \dfrac{F_g - W}{\rho_w g}\\
V = \dfrac{96N - 32N}{1000kg/m^3\cdot 10N/kg} = 6.4\times 10^{-3}m^3"

Answer. "6.4\times 10^{-3}m^3".